OpenStudy (anonymous):

Select the best choice for the value f(a-2) of the dependent variable for the function f(x)=-2x^2-3x-1. A.) y=-5a+9 B.) y=2a-3 C.) y=2a^2-11a-13 D.) y=-2a^2+5a-3 I think I kinda get the idea, but I dont understand the question. Its the beginning of the year and I plan on acing this class so I want to understand EVERYTHING.

OpenStudy (anonymous):

okay so dependent variable is y, or f

OpenStudy (anonymous):

so the value for the variable x is a-2. So all they want to know is what is the value of y when x=a-2 for the equation f)x)=-2x^2-3x-1.

OpenStudy (anonymous):

so, y=-2(a-2)^2-3(a-2)-1

OpenStudy (cwrw238):

yes plug in x = a-2 in f(x)

OpenStudy (anonymous):

so f=y right?

OpenStudy (cwrw238):

yes

OpenStudy (cwrw238):

simplify and see what you get

OpenStudy (anonymous):

Alrighty. They just have to be complicated. :) y=-2(a-2)^2-3(a-2)-1 y=-2(a^2-4)-3a-6-1 y=-2a^2-8-3a-7 y=2a^2-3a-15

OpenStudy (anonymous):

>:/ which isn't an option.

OpenStudy (anonymous):

you made a mistake when calculating (a-2)^2

OpenStudy (anonymous):

hm actually in another place too, a negative number multiplied by another negative number equal a positive number

OpenStudy (cwrw238):

yes (a-2)^2 = a^2 -4x + 4

OpenStudy (anonymous):

yes (a-2)^2 = a^2 -4x + 4 not understanding what that is...and it'd be a^2 and then 2squared is 4, so a^2-4.

OpenStudy (anonymous):

have you learned to FOIL yet? (a-2)^2 doesnt equal (a^2) +(-2)^2, its (a-2)(a-2)

OpenStudy (anonymous):

FOIL? I've never heard of that.

OpenStudy (cwrw238):

= -2(a^2 -4a + 4) -3a + 5 = -2a^2 +8a - 8 - 3a+ 5 # -2a^2 + 5a - 3

OpenStudy (anonymous):

Where did you get the 4a +4?

OpenStudy (anonymous):

I promise I'm not trying to be stupid :)

OpenStudy (anonymous):

ok while he explains that, I just want to clarify that -3(a-2) = -3(a)+(-3)(-2) = -3a +6

OpenStudy (cwrw238):

(a - 2)^2 = (a - 2)(a - 2) now multiply the second bracket by a then by -2 = a(a - 2) +2(a - 2) = a^2 - 2a - 2a + 4 = a^2 - 4a + 4

OpenStudy (cwrw238):

second line should read = a(a - 2) - 2(a - 2

OpenStudy (anonymous):

alternatively you could use http://en.wikipedia.org/wiki/FOIL_method so (a-2)(a-2) = a^2 -2a -2a +4 = a^2 -4a +4

OpenStudy (anonymous):

oh okay...so let me start over then. y=-2(a-2)^2-3(a-2)-1 y=-2[(a-2)(a-2)]-3(a-2)-1 y=-2(a^2-2a-2a+4)-3a+6-1 y=-2a^a-3a+9

OpenStudy (anonymous):

last line should be y=-2a^2-7a+9

OpenStudy (anonymous):

*facepalm*

OpenStudy (cwrw238):

you've gone wrong on last line next step is to multiply the terms in the parentheses by -2

OpenStudy (cwrw238):

dont try and do too much processing at same time

OpenStudy (anonymous):

yes yes i see that... okay AGAIN. :) y=-2(a^2-2a-2a+4)-3a+6-1 SO, y=-2a^2+8a-8-3a+6-1 therefore... y=-2a^2+5a-3

OpenStudy (anonymous):

Which is an option hurrah!

OpenStudy (cwrw238):

yea - you got the right option too

OpenStudy (anonymous):

:) thank you so much, that (a-2)(a-2) helped

OpenStudy (cwrw238):

yw - you can use FOIL if you like personally i prefer the other method where you multiply the second bracket by each part of first bracket - thats call distributing a and -2 over the (x - 2) - The distributive law

OpenStudy (anonymous):

Okay, I'll go the long a sure route, thats the way my brain works. :)